# Towards a Robust Transport Network With Self-adaptive Network Digital Twin

### v1

@@ 3 Uniform 5G-Crosshaul 347 002347\,002347 002
The use case scenario used to evaluate the proposed NDT synchronization, considering a high-level metric related to SLA monitoring tasks, was carried out using only data coming from 5G-crosshaul topology. In this sense, we evaluated the same scenario presented in Figure 8, but considering high-level metrics such as the accuracy in the SLA violation predictions. Thus, we assume that each flow has a specific PDB value that was assigned empirically, based on the characteristics of the 5G-crosshaul topology, such as the propagation delay of each link.
Regarding the experiments, we consider a total of 25 10025\,10025 100 flows traversing the network, each of which must be classified by the SLA monitoring application as compliant or in violation with the PDB, based on the delay predicted by the VTwin. The SLA violation in this case is characterized when a per-flow delay exceeds the PDB for a given path. From the total of flows considered in this experiment, 536536536 flows are violating with their SLA (ground truth). Hence, the accuracy in the prediction of SLA violations is shown in Figure 10, which considers the VTwin with the NDT synchronization module and without it, and each window index corresponds to the classification of 100 flows, for a better visualization of the results in this figure.
Figure 10: SLA violations in scenario without (bottom) and with twin synchronization (upper). The orange curve show the number of actual SLA validation, and the blue curve the number of predicted violations.
@@
As shown in Figure 10, the application exhibits the expected degradation in classification performance when the NDT synchronization module is disabled, due to the VTwin states not being updated with PTwin states. Quantitatively, this results in a misclassification of 7 7447\,7447 744 flows that are not in compliance with their respective SLAs. In contrast, when the NDT synchronization module is enabled, the number of incorrectly classified flows is significantly reduced to 3 $3$\,$3$ 333. It is important to note that part of these SLA violations also occurs during periods when the VTwin is undergoing retraining. Overall, this experiment highlights that the practical usability of an NDT strongly depends on the presence of an effective synchronization mechanism between the twins, operating at an acceptable rate, which is essential for transitioning the solution from a prototype to a production-ready platform or framework.
@@ 5 Conclusions and Future Works
[…]
@@
The results of the proposed experiments offer valuable insights into the challenges of twin synchronization, positioning this article as a foundational step toward future developments. A key direction for future work involves progressively incorporating more realistic elements into the experimental environments. One major limitation is the gap between the current setup and real-world production environments. Although the proposed module presents a promising approach to addressing specific challenges faced by NDT systems, it does not fully resolve all the issues required for the complete feasibility of the NDT platform. For instance, while the data management module includes essential mechanisms for collecting information from the PTwin, the experiments rely on traffic generated by a traffic generator, with metrics directly obtained from simulator logs. In contrast, data collection in a production environment is significantly more complex and may require integration with auxiliary tools, such as those defined by protocols such as the Simple Network Management Protocol (SNMP) used to collect network metrics. Furthermore, the feasibility of the proposed synchronization method also depends on the ability to generate synthetic data to retrain the VTwin model when necessary, which is not addressed in this article. In this context, the use of generative AI emerges as a promising solution and represents a potential avenue for future research.
@@
In addition to the data management tasks associated with an NDT, future research will also focus on improving the synchronization method itself, which can be pursued along two main directions. In the short term, the goal is to evaluate synchronization performance in variable transport network environments using a more robust multivariate concept drift detection approach, one that simultaneously considers multiple input features such as traffic rate, per-link load, and link capacity. In the long term, we aim to develop a more proactive synchronization strategy by integrating online learning techniques. This would reduce the latency in adapting the VTwin model parameters without waiting for an explicit concept drift detection, thereby minimizing the duration which outdated model weights are in effect.
@@ Acknowledgments
Data Availability
Data will be made available on request.

### v2

@@ 4 Deterministic 5G-Crosshaul 100 800100\,800
The use case scenario employed to evaluate the proposed NDT synchronization, considering a high-level metric related to SLA monitoring tasks, was conducted using data derived from the four topologies used in the last experiments. In this sense, we evaluated the same scenario used to validate the performance on prediction per-flow delay, but we considered high-level metrics such as the accuracy of the SLA violation predictions. In this case, we assume a simplified SLA compliance scenario, admitting only the predicted per-flow delay as the metric used to make a decision regarding the SLA violation. Thus, we assume that each flow has a specific PDB value that was assigned empirically based on the characteristics of each topology, such as the propagation delay of each link.
Regarding the experiments, we considered a total of 25 10025\,100, 24 60024\,600, 26 70026\,700, and 39 70039\,700 flows traversing the 5G-Crosshaul, Germany, PASSSION, and Synthetic-700 network topologies, respectively. Each flow was evaluated by the SLA monitoring application to determine whether it was compliant with or in violation of the PDB, based on the delay predicted by the VTwin. An SLA violation occurs when the per-flow delay exceeds the PDB for the corresponding path. Table 9 summarizes the number of flows identified as being in violation for each topology.
Table 9: Flow traffic characteristics for each topology used for SLA compliance classification.
Topology Total number of flows Number of flows in violation
5G-Crosshaul 25 10025\,100 $536$
Germany 24 60024\,600 $793$
PASSION 26 70026\,700 1 2291\,229
Synthetic-700 39 70039\,700 1 5271\,527
The accuracy in the prediction of SLA violations is shown in Figure 19(a) and 19(b), which considers the VTwin with and without the NDT synchronization module in the 5G-Crosshaul and Germany topologies. Each window index corresponds to the classification of 100 flows for better visualization of the results in this figure. In all topologies, the application exhibits the expected degradation in classification performance when the NDT synchronization module is disabled. This occurs because the VTwin states are not updated with those of the PTwin. Quantitatively, this results in 8 1158\,115 and 6 3476\,347 flows being incorrectly classified as compliant with their respective SLAs, in the 5G-Crosshaul and Germany topologies, respectively.
(a) Accuracy of SLA violation predictions in 5G-Crosshaul topology.
(b) Accuracy of SLA violation predictions in Germany topology.
Figure 19: Accuracy of SLA violation classification considering 5G-Crosshaul and Germany topologies.
@@
When the NDT synchronization module is enabled, however, the number of misclassified flows decreases significantly to 4 5364\,536 and $982$, respectively. Similar improvements are observed in the other two topologies, as depicted in Figure 20(a) and 20(b), respectively, the PASSION and Synthetic-700 topologies. Moreover, all numerical results are summarized in Table 10. It is also noteworthy that part of these SLA violations occurs during periods in which the VTwin is undergoing retraining. Overall, this experiment demonstrates that the practical usability of an NDT critically depends on an effective synchronization mechanism between the twins operating at an appropriate rate, which is essential for transitioning the solution from a prototype to a production-ready framework.
(a) Accuracy of SLA violation predictions in PASSION topology.
(b) Accuracy of SLA violation predictions in Synthetic-700 topology.
Figure 20: Accuracy of SLA violation classification considering PASSION and Synthetic-700 topologies.
Table 10: Application performance on SLA violation classification with and without NDT synchronization.
Without NDT synchronization With NDT synchronization
Topology Number of flows misclassified Number of flows correctly classified Accuracy (%) Number of flows misclassified Number of flows correctly classified Accuracy (%)
5G-Crosshaul 8 1158\,115 16 98516\,985 $67.66$ $\mathbf{4\,536}$ $\mathbf{20\,564}$ $\mathbf{81.92}$
Germany 6 3476\,347 18 25318\,253 $74.19$ $\mathbf{982}$ $\mathbf{23\,618}$ $\mathbf{96.00}$
PASSION 9 0199\,019 17 68117\,681 $66.22$ $\mathbf{3\,941}$ $\mathbf{22\,759}$ $\mathbf{85.23}$
Synthetic-700 12 13312\,133 27 59727\,597 $69.43$ $\mathbf{7\,788}$ $\mathbf{31\,912}$ $\mathbf{80.38}$
@@ 5 Conclusions and Future Works
This article proposed an enhanced NDT architecture incorporating data-driven synchronization modules, taking initial steps toward addressing practical and production-level aspects of NDT deployment in transport networks. The proposed modules within this architecture specifically tackle the challenge posed by the non-stationary nature of network traffic over time. In this sense, this architecture aims to ensure synchronization between the two twins of the NDT platform. To achieve this, the NDT synchronization modules, such as the concept drift detector, aim to track and detect possible changes in the traffic pattern, triggering a retraining process and ultimately updating the VTwin weights through the VTwin model management. To evaluate the architecture’s operation, we considered four different topologies with sizes ranging from 50 to 700 nodes in an environment with flows using four packet size distributions, generating various traffic patterns. Thus, to detect and initiate the retraining process, we considered an input distribution-based approach, which was based on the KSWIN method. Our results demonstrate the importance of the synchronization process between the VTwin and PTwin, achieving a performance improvement in predicting total per-flow and per-flow jitter of at least $64\%$ and $21\%$, respectively, compared to a configuration without a concept drift detector. Finally, aimed at obtaining insight into the impact of the proposed NDT architecture at a high level, we also evaluated the NDT synchronization, considering an application related to SLA monitoring.
@@
The results of the proposed experiments offer valuable insights into the challenges of twin synchronization, positioning this article as a foundational step toward future developments. A key direction for future work involves progressively incorporating more realistic elements into the experimental environments. One major limitation is the gap between the current setup and real-world production environments. Although the proposed module presents a promising approach to addressing specific challenges faced by NDT systems, it does not fully resolve all the issues required for the complete feasibility of the NDT platform. For instance, while the data management module includes essential mechanisms for collecting information from the PTwin, the experiments rely on traffic generated by a traffic generator, with metrics directly obtained from simulator logs. In contrast, data collection in a production environment is significantly more complex and may require integration with auxiliary tools, such as those defined by protocols such as the Simple Network Management Protocol (SNMP), which is used to collect network metrics. Furthermore, the feasibility of the proposed synchronization method also depends on the ability to generate synthetic data to retrain the VTwin model when necessary, which is not addressed in this article. In this context, the use of generative AI emerges as a promising solution and represents a potential avenue for future research.
@@
In addition to addressing the data management tasks inherent to an NDT, future research will also focus on enhancing the synchronization method itself, which can be explored along two main directions. In the short term, we aim to assess synchronization performance under varying transport network conditions by adopting a more robust multivariate concept drift detection approach, one that simultaneously accounts for multiple input features, such as traffic rate, per-link load, and link capacity. Moreover, in future work, we will also consider incorporating a more realistic distribution of packet rates, as the current study assumes the same distribution for packet rate and packet size. Another aspect that should be evaluated in future work is the performance impact of the VTwin retraining process caused by false positives from these concept drift detectors, which is not addressed in this present work. In the long term, we aim to develop a more proactive synchronization strategy by integrating online learning techniques. This approach would reduce the latency in adapting the VTwin model parameters by eliminating the need to wait for explicit concept drift detection, thereby minimizing the period during which outdated model weights remain in use.